Solve for y
y = 2(-1/5) + 9

Answer by jim_thompson5910(35256)

y = 2(-1/5) + 9


y = (2/1)(-1/5) + 9


y = (2*-1)/(1*5) + 9


y = -2/5 + 9


y = -2/5 + (9/1)*(5/5)


y = -2/5 + (9*5)/(1*5)


y = -2/5 + 45/5


y = (-2 + 45)/5


y = 43/5


-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------